A laser produces a beam of light with a wavelength of 532 nm. If the power output is 25.0 mW, how many photons does the laser emit per second? (W = 1 J/s)
1
Expert's answer
2020-11-22T11:12:09-0500

The energy of a photon (E) = "\dfrac{hc}{\lambda}"


E = "\dfrac{6.63×10^{-34}×3×10^8}{532× 10^{-9}}"


"= 3.74× 10^{-19}\ J"


Power output = 25mW = "25 \u00d7 10^{-3}"W.

This means, the energy emitted in one second = "25 \u00d7 10^{-3}"J


Number of photons in one second = "\dfrac{25×10^{-3}}{3.74×10^{-19}} = 6.68 × 10^{16}"